The firm has an average collection period of 34 days. Current practice is to factor all receivables immediately at a discount rate of (1+0.1*x)%. Assume that default is extremely unlikely. What is the effective annual interest rate on this arrangement? Let x=1

A) 11.39 percent

B) 12.61 percent

C) 13.84 percent

D) 15.08 percent

E) 16.34 percent
Answer #1

Answer is 12.61%

Discount Rate = (1 + 0.1*x)%
Discount Rate = (1 + 0.1*1)%
Discount Rate = 1.10%

Average Collection Period = 34 days

Effective Interest Rate = [1 + Discount Rate / (1 - Discount Rate)]^[365 / Average Collection Period] - 1
Effective Interest Rate = [1 + 0.0110 / (1 - 0.0110)]^[365 / 34] - 1
Effective Interest Rate = [1 + 0.0110 / 0.9890]^[365 / 34] - 1
Effective Interest Rate = 1.011122^10.735294 - 1
Effective Interest Rate = 1.1261 - 1
Effective Interest Rate = 0.1261 or 12.61%